OpenStudy (anonymous):

how do you solve: 7/4x+7/12=1/3x-1/6. Please show steps.

OpenStudy (anonymous):

\[\frac{ 7 }{4 }x+\frac{ 7 }{12 }=\frac{ 1 }{3 }x-\frac{ 1 }{6 }\]

OpenStudy (anonymous):

Anyone know how to solve this?